OpenStudy (anonymous):

A can contains 14 pennies, 12 nickles, 16 dimes, and 8 quarters. What is the probability of drawing either a penny or a nickle on one draw?

OpenStudy (kropot72):

The events 'draw a penny' and 'draw a nickle' are mutually exclusive. Therefore we can find the probability of drawing either a penny or a nickle on one draw by adding the probabilities of 'draw a penny' and 'draw a nickle'. \[P(draw\ a\ penny)=\frac{14}{50}\] \[P(draw\ a\ nickle)=\frac{12}{50}\] Can you finish the calculation now?

OpenStudy (anonymous):

@_Michaeld22

OpenStudy (kropot72):

\[\frac{14}{50}+\frac{12}{50}=\ ?\]

OpenStudy (anonymous):

13/25

OpenStudy (kropot72):

You are correct!

OpenStudy (anonymous):

thank u

OpenStudy (kropot72):

You're welcome :)